### v1

# PoLi-RL: A Point-to-List Reinforcement Learning Framework for Conditional Semantic Textual Similarity

Abstract
Conditional Semantic Textual Similarity (C-STS) measures the semantic proximity between text segments under a specific condition, thereby overcoming the ambiguity inherent in traditional STS. However, existing methods are largely confined to discriminative models, failing to fully integrate recent breakthroughs in the NLP community concerning Large Language Models (LLMs) and Reinforcement Learning (RL). RL is a particularly well-suited paradigm for this task, as it can directly optimize the non-differentiable Spearman ranking metric and guide the reasoning process required by C-STS. However, we find that naively applying listwise RL fails to produce meaningful improvements, as the model is overwhelmed by complex, coarse-grained reward signals. To address this challenge, we introduce PoLi-RL, a novel Point-to-List Reinforcement Learning framework. PoLi-RL employs a two-stage curriculum: it first trains the model with simple pointwise rewards to establish fundamental scoring capabilities, then transitions to a hybrid reward that combines pointwise, pairwise, and listwise objectives to refine the model’s ability to discern subtle semantic distinctions. Crucially, we propose an innovative Parallel Slice Ranking Reward (PSRR) mechanism that computes ranking rewards in parallel slices, where each slice comprises same-indexed completions from different samples. This provides a precise, differentiated learning signal for each individual completion, enabling granular credit assignment and effective optimization. On the official C-STS benchmark, PoLi-RL achieves a Spearman correlation coefficient of 48.18, establishing a new SOTA for the cross-encoder architecture. As the first work to successfully apply RL to C-STS, our study introduces a powerful and precise paradigm for training LLMs on complex, ranking-based conditional judgment tasks.
11footnotetext: Equal Contribution. Work done during the internship at Tencent Youtu Lab.
1 Introduction
As a core research area in Computational Linguistics, Semantic Textual Similarity (STS) (Agirre et al., 2013) finds extensive applications across diverse scenarios, including topic modeling, dialogue systems, text summarization, and agent memory (Tang et al., 2025). However, traditional STS tasks exhibit inherent ambiguity because similarity definitions are often susceptible to observer bias. To address this limitation, the Conditional Semantic Textual Similarity (C-STS) task was developed (Deshpande et al., 2023). By incorporating an explicit natural language condition, C-STS enables more precise and objective similarity judgments, yet simultaneously imposes higher demands on a model’s reasoning capabilities. For instance, consider the following two text fragments: “A player is shooting from beyond the three-point line” and “A player is taking a free throw”. Under the condition “The activity of the player”, their similarity is high. However, under the condition “The player’s distance from the basket”, their similarity is low.
Research on this nascent task has yielded three primary paradigms: Bi-encoder (Liu et al., 2025), Tri-encoder (Lin et al., 2024), and Cross-encoder (Li et al., 2024). The Cross-encoder architecture, which processes the text pair and the guiding condition simultaneously, is the most compatible with modern generative pre-trained models. Despite this, the integration of C-STS with LLMs remains in its early stages. Current LLM applications are limited to two main approaches: direct inference via few-shot prompting, where even state-of-the-art models struggle to achieve satisfactory results (Deshpande et al., 2023); and their use as feature extractors for generating text embeddings (Yamada and Zhang, 2025), which is an extension of the discriminative paradigm. To the best of our knowledge, no prior work has applied an end-to-end LLM-based cross-encoder to the C-STS task, nor has any integrated it with advanced training techniques like reinforcement learning (RL), leaving a significant research gap.
This paper aims to fill this gap. We posit that incorporating RL into an LLM-based cross-encoder paradigm is a natural fit. This is reflected in two aspects: First, C-STS requires nuanced, scenario-based reasoning. For example, in the basketball case described earlier, to correctly assess similarity under the “distance” condition, the model must move beyond surface-level semantics to identify the underlying spatial relationship between ‘beyond the three-point line’ and ‘at the free throw line’, a process demanding strong abstraction and inference. RL, through its explicit reward signals, can more effectively guide the reasoning process of LLMs (Guo et al., 2025). Second, from an optimization standpoint, RL aligns closely with the task’s evaluation criteria. The Spearman correlation coefficient (Zar, 2005), a core evaluation metric of C-STS, is a non-differentiable measure of ranking quality. Traditional SFT methods can only indirectly and approximately optimize this objective through loss functions like Mean Squared Error (MSE) (Zhang and Li, 2024). In contrast, RL allows for the direct optimization of ranking-based reward functions that are designed to correlate strongly with the final Spearman metric, maintaining a higher degree of consistency with the final evaluation.
However, a naive application of RL to this task presents significant challenges. As illustrated in Figure 1, our preliminary experiments indicate that directly applying a single listwise ranking reward (e.g., Spearman’s correlation coefficient) across an entire batch of completions does not show any advantages compared to the few-shot method. This approach suffers from two fundamental problems. First, the ranking objective is too complex for a model that has not yet learned the task’s fundamental scoring rules, often causing the training to stagnation. Second, a single reward computed across the entire batch is too coarse to provide precise credit assignment, as a few poor completions can unfairly penalize other good ones.
Figure 1: Performance comparison of different strategies on the C-STS task. Directly applying listwise ranking rewards for RL does not significantly outperform the few-shot baseline. In contrast, both stages of our method (PoLi-RL) achieve substantial improvements, validating its effectiveness.
To address these challenges, we propose PoLi-RL, a two-stage Point-to-List Reinforcement Learning framework. PoLi-RL features a two-stage curriculum to manage the complexity of the learning task. In the first stage, we use simple pointwise rewards to ground the model in the basic scoring rules of the task. Building on this foundation, the second stage introduces a richer, hybrid reward signal that combines a stable pointwise anchor with more nuanced pairwise and listwise ranking rewards, enabeling the model to discern subtle semantic differences.
Furthermore, to resolve the problem of coarse-grained reward signals that arises from ranking all completions in a batch together, we innovatively introduce a Parallel Slice Ranking Reward (PSRR) mechanism, which utilizes a two-level decomposition. First, for a batch of input samples, the model generates $G$ completions for each. We then form $G$ “parallel slices”, where the $i$-th slice is composed of the $i$-th completion from every sample. Second, and more importantly, within each slice, rather than assigning a single reward, we compute the rank difference for each individual completion against its ideal rank. This two-level decomposition allows each of the $N\times G$ completions to receive a unique and precise reward that reflects its quality, thereby facilitating granular credit assignment and more effective optimization.
The primary contributions of this paper are outlined as follows:
  • To the best of our knowledge, this is the first work to propose an end-to-end, LLM-based cross-encoder for the C-STS task and the first to employ reinforcement learning for training in this domain.
  • We design and implement PoLi-RL [^1] The code is available at: https://github.com/ZBWpro/PoLi-RL., a novel two-stage training curriculum that addresses the complexity of C-STS by progressing from a simple pointwise reward to a more complex hybrid reward.
  • We propose the Parallel Slice Ranking Reward (PSRR) mechanism, which delivers precise and differentiated learning signals by computing ranking rewards within independent ‘parallel slices’ to enable granular credit assignment. This mechanism offers a generalizable strategy for other ranking and retrieval tasks involving multiple generation candidates.
  • On the official C-STS benchmark, PoLi-RL achieves a Spearman’s correlation coefficient of 48.18, establishing a new SOTA for the cross-encoder architecture and surpassing strong closed-source models, including GPT-4 (43.6) (Achiam et al., 2023). Our qualitative analysis further reveals our method’s advantages in understanding complex conditions.
2 Methodology
This section details our proposed strategy. We begin by formulating the C-STS task within an end-to-end, LLM-based cross-encoder paradigm in subsection 2.1. Then, in subsection 2.2, we map this task onto the mathematical framework of Reinforcement Learning and specifiy its optimization objective. Finally, in subsection 2.3, we provide a comprehensive description of our PoLi-RL framework, including its two-stage design and the innovative PSRR mechanism.
2.1 Problem Formulation
The core objective of C-STS is to learn a scoring function that accurately reflects the semantic similarity between two text segments under a specific condition. Formally, each C-STS data sample is defined as a tuple $x=(t_{1},t_{2},c,y)$, where $t_{1},t_{2}$ are two text segments, $c$ is the natural language condition, and $y\in[1,5]$ is the human-annotated similarity judgement on the Likert scale (Likert, 1932). Notably, the label $y$ corresponds to a fine-grained set of semantic criteria. According to the C-STS annotation guidelines, the meanings of the scores are as follows: (1) Completely dissimilar; (2) Thematically related but dissimilar; (3) Roughly equivalent, but with some important information differences; (4) Mostly equivalent, with some unimportant details differing; (5) Completely equivalent. This level of granularity demands that the model perform fine-grained reasoning beyond surface-level semantics, posing a significant challenge to its capabilities.
A unique characteristic of the C-STS dataset is its paired structure: samples are organized in adjacent pairs that share the same text segments $(t_{1},t_{2})$ but feature different conditions and maintain a deterministic ordinal relationship between their labels, i.e., $y_{\text{high}}\geq y_{\text{low}}$. This structure provides a solid foundation for our pairwise reward design, as detailed in subsection 2.3.
Our task is to train a scoring model $\pi_{\theta}$, parameterized by $\theta$. For each sample $x$, the model takes a unified prompt $p=\text{INS}(x)$ as input, which is constructed from the text segments, condition, and task-specific instructions (detailed prompt templates are provided in A.2), to generate an output sequence $o=\pi_{\theta}(p)$. From this sequence, we parse the final predicted score, $\tilde{y}=\text{Parse}(o)$. The overall training objective is to maximize the ranking consistency between the set of predicted scores $\{\tilde{y}_{i}\}_{i=1}^{N}$ and the ground-truth scores $\{y_{i}\}_{i=1}^{N}$ by optimizing the policy $\pi_{\theta}$, a process primarily measured by Spearman’s correlation coefficient. Since this metric is rank-based and non-differentiable, RL emerges as a more promising optimization paradigm than traditional supervised fine-tuning.
2.2 Reinforcement Learning for C-STS
We formulate the C-STS task as a Markov Decision Process (MDP), defined by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma)$, where the agent is the LLM policy $\pi_{\theta}$. The generation process is modeled as a sequence of decisions, where each step involves generating a single token. A state $s_{t}\in\mathcal{S}$ at timestep $t$ is the sequence of tokens generated so far, conditioned on the initial prompt, i.e., $s_{t}=(p,o_{<t})$. An action $a_{t}\in\mathcal{A}$ is the selection of the next token $o_{t}$ from the model’s vocabulary, governed by the policy $\pi_{\theta}(a_{t}|s_{t})$, which provides a probability distribution over all possible tokens. The transition function $\mathcal{T}$ is deterministic given a selected action: the next state $s_{t+1}$ is formed by appending the chosen token $a_{t}$ to the current state $s_{t}$. The stochasticity in the generation process arises from the sampling procedure, where actions are drawn from the policy distribution $\pi_{\theta}(a_{t}|s_{t})$ with configurable parameters such as temperature and top-p. We employ a terminal reward setting, where a reward $\mathcal{R}_{T}=\mathcal{R}(x,o)$ is given only after the entire sequence $o$ has been generated. Finally, the discount factor $\gamma$ is set to 1 to ensure that the terminal reward is backpropagated without decay to all actions that contributed to the final output. Based on this framework, the objective is to find the optimal parameters $\theta^{*}$ that maximize the expected reward over the data distribution $\mathcal{D}$:
$\theta^{*}=\underset{\theta}{\operatorname{arg\,max}}\mathbb{E}_{x\sim\mathcal{D},o\sim\pi_{\theta}(p)}[R(x,o)]$ (1)
To optimize this objective, we employ Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) (Yu et al., 2025), an extension of GRPO (Shao et al., 2024) that introduces several key techniques for effective RL. For each sample $x$, the policy generates a set of $G$ completions $\{o_{i}\}_{i=1}^{G}$. A scalar reward $r_{i}=R(x,o_{i})$ is computed for each completion. The advantage $\hat{A}_{i}$ for each completion is then calculated by normalizing its reward against the statistics of the entire group’s rewards via Z-score normalization:
$\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r_{i}\}_{i=1}^{G})}{\text{std}(\{r_{i}\}_{i=1}^{G})+\epsilon}$ (2)
Since we adopt an on-policy training approach, completions are sampled directly from the current policy $\pi_{\theta}$ being optimized. The objective function for updating the model parameters $\theta$ is:
$\mathcal{J}_{\text{DAPO}}(\theta)=\mathbb{E}_{x\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta}(\cdot|p)}\left[\frac{1}{\sum_{i=1}^{G}|o_{i}|}\sum_{i=1}^{G}\sum_{t=1}^{|o_{i}|}\left(\frac{\pi_{\theta}(o_{i,t}|p,o_{i,<t})}{[\pi_{\theta}(o_{i,t}|p,o_{i,<t})]_{\text{nograd}}}\hat{A}_{i,t}\right)\right]$ (3)
where $\cdot]_{\text{nograd}}$ denotes stopping the gradient flow through the denominator, ensuring that only the numerator is updated during backpropagation.
2.3 PoLi-RL: A Two-Stage Reinforcement Learning Framework
As previously established, our optimization goal is to maximize the expected reward, making the design of the reward function $\mathcal{R}$ central to our method. To ensure stable and effective optimization, we propose PoLi-RL, a framework built upon a two-stage progressive reward curriculum. This subsection details our pipeline and its reward mechanisms, as illustrated in Figure 2.
Figure 2: An overview of the PoLi-RL framework. The framework employs a two-stage curriculum, progressing from Stage I, where the model learns foundational scoring rules, to Stage II, which refines the model’s ability to discern fine-grained semantic differences. The core of our method is the PSRR mechanism in Stage II, where pairwise and listwise ranking rewards are computed vertically within slices of same-indexed completions to provide precise, differentiated learning signals.
Stage I: Foundational Skill Acquisition.
The goal of Stage I is to ground the model in the fundamental scoring rules of the C-STS task. For each input sample, the policy generates $G$ completions, from which we parse a set of predicted scores $\{\tilde{y}_{j}\}_{j=1}^{G}$. The total reward for Stage I, $R_{S1}$, is a weighted sum of three components:
$R_{S1}=\lambda_{1}R_{\text{pointwise}}+\lambda_{2}R_{\text{binary}}+\lambda_{3}R_{\text{format}}$ (4)
The Pointwise Accuracy Reward ($R_{\text{pointwise}}$) is the primary component in Stage I. It measures the normalized distance between the predicted score $\tilde{y}_{j}$ and the ground-truth score $y_{j}$.
$R_{\text{pointwise}}=1-\frac{|\tilde{y}_{j}-y_{j}|}{\max(Y)-\min(Y)}$ (5)
where $\max(Y)=5$ and $\min(Y)=1$ are the bounds of the label space.
To counter reward hacking, where the model tends to output safe intermediate scores, we introduce a Binary Judgement Reward ($R_{\text{binary}}$). According to the C-STS guideline that scores $\geq 3$ indicate similarity while scores $\leq 2$ indicate dissimilarity, this reward encourages the model to first master this basic binary classification:
$R_{\text{binary}}=\begin{cases}1&\text{if }(\tilde{y}_{j}\geq 3\land y_{j}\geq 3)\lor(\tilde{y}_{j}<3\land y_{j}<3)\\ 0&\text{otherwise}\end{cases}$ (6)
Finally, a simple Format Reward ($R_{\text{format}}$) ensures the output adheres to the required structure, which consists of a binary judgment (‘yes’ or ‘no’) followed by the numerical score in parentheses.
Stage II: Fine-Grained Semantic Refinement.
After the model has acquired basic scoring abilities in Stage I, Stage II refines its capacity to discern subtle semantic differences by introducing a richer, hybrid reward signal that incorporates both pairwise and listwise ranking objectives.
PSRR: A Fine-Grained Reward Mechanism.
A primary challenge in directly optimizing ranking metrics is that a single, batch-wide reward signal is too coarse to assign credit precisely. To address this, we propose the Parallel Slice Ranking Reward (PSRR) mechanism. The core idea of PSRR is to restructure the generated outputs to enable more granular reward computation. For a batch of $N$ samples, we begin by having the policy generate $G$ completions $\{o_{i,1},\dots,o_{i,G}\}$ for each sample $x_{i}$. From each completion $o_{i,j}$, we then parse the predicted score $\tilde{y}_{i,j}$. Instead of treating the completions as a single flat list, we organize these $N\times G$ predicted scores into $G$ “parallel slices”. Each slice, denoted as $Y^{j}_{pred}$, is defined as the collection of the $j$-th predicted score from all $N$ samples in the batch: $Y^{j}_{pred}=\{\tilde{y}_{1,j},\tilde{y}_{2,j},\ldots,\tilde{y}_{N,j}\}$, where $j\in\{1,\ldots,G\}$. This slicing architecture is the foundation upon which our advanced ranking rewards are built, ensuring that each completion receives a specific learning signal based on its relative performance within its slice.
A sufficiently large slice size $N$ is crucial for a stable and meaningful ranking signal. To make this computationally feasible with limited GPU memory, we leverage gradient accumulation. Specifically, we first generate the full set of $N\times G$ completions across multiple devices and accumulation steps. These completions are then organized into parallel slices to compute rewards and advantages globally. Subsequently, for the forward and backward passes that compute loss and gradients, we process smaller sub-batches sequentially, accumulating gradients over multiple passes before executing a single optimizer step. This strategy makes our reward design practically feasible, enabling the model to learn from a rich, large-scale ranking signal without requiring prohibitive memory.
Pairwise Ranking Reward.
Computed within each parallel slice $P_{j}$, this reward leverages the paired structure of the C-STS dataset to provide a local ranking signal. It is applied only to adjacent input samples ($x_{i},x_{i+1}$) that form a valid pair. For such a pair, we define the predicted difference as $\Delta_{\text{pred}}=\tilde{y}_{i,j}-\tilde{y}_{i+1,j}$ and the true difference as $\Delta_{\text{true}}=y_{i}-y_{i+1}$. The reward $R^{\text{pairwise}}_{i,j}$ is then a piecewise function:
$R^{\text{pairwise}}_{i,j}=\begin{cases}0&\text{if }\text{sgn}(\Delta_{\text{pred}})\neq\text{sgn}(\Delta_{\text{true}})\\ R_{\text{base}}+(1-R_{\text{base}})\cdot\left(1-\frac{\left|\Delta_{\text{pred}}-\Delta_{\text{true}}\right|}{\text{max\_error}}\right)&\text{if }\text{sgn}(\Delta_{\text{pred}})=\text{sgn}(\Delta_{\text{true}})\end{cases}$ (7)
This function first checks if the basic ranking preference is correct using the sign function $\text{sgn}(\cdot)$. If the order is wrong, the reward is zero. If correct, a base reward $R_{\text{base}}$ is given, plus an additional reward that measures the normalized distance between the predicted and true score differences. Here, max_error is the maximum possible score difference, which is 3 for paired samples.
Listwise Ranking Reward.
While the pairwise reward focuses on local comparisons, the listwise reward provides a more global ranking perspective within each slice. It is calculated as the normalized difference between a completion’s predicted rank within its slice and the ideal rank of its ground-truth label, formulated as:
$R^{\text{listwise}}_{i,j}=1-\frac{|\text{Rank}(\tilde{y}_{i,j},Y^{j}_{pred})-\text{Rank}(y_{i},Y_{\text{true}})|}{N-1}$ (8)
where $Y_{\text{true}}=\{y_{1},\ldots,y_{N}\}$ is the set of true labels for the current batch, the function $\text{Rank}(v,S)$ returns the rank of element $v$ within the set $S$ (from 1 to $N$), and the division by N1N-1 normalizes the rank error to the range $[0,1]$.
The final reward for Stage II, $R_{S2}$, combines the robust Pointwise Reward from Stage I as a stabilizing anchor with the new ranking-based rewards enabled by PSRR. The total reward is a weighted combination of these three components:
$R_{S2}=\mu_{1}R_{\text{pointwise}}+\mu_{2}R_{\text{pairwise}}+\mu_{3}R_{\text{listwise}}$ (9)
3 Experiments
To empirically validate the effectiveness of our proposed PoLi-RL framework, we conduct a comprehensive set of experiments. We begin by detailing our experimental setup in subsection 3.1, including the dataset, evaluation metrics, baselines, and implementation details. Following this, in subsection 3.2, we present the main results, comparing PoLi-RL against a suite of strong baselines. Finally, in subsection 3.3, we conduct a series of ablation studies to analyze the contributions of our framework’s key components.
3.1 Experimental Setup
We build PoLi-RL upon the Qwen3-8B (Yang et al., 2025) model using the ms-swift framework (Zhao et al., 2024) for RL training. All experiments are conducted on the official C-STS dataset (Deshpande et al., 2023). Following prior work, we use Spearman correlation as the primary metric and Pearson as the secondary. We compare our method against three baseline categories: First, the discriminative models in the cross-encoder setting. Second, powerful generative LLMs, such as Flan-UL (Tay et al., 2022), Flan-T5 (Chung et al., 2024) and Tk-Instruct (Wang et al., 2022). Finally, our own SFT and few-shot implementations on Qwen3-8B for direct comparison.
Table 1: Main results on the official C-STS benchmark. All scores are reported as Spearman/Pearson correlation coefficients multiplied by 100. Results marked with $\dagger$ are obtained from (Deshpande et al., 2023), while $\ddagger$ denotes results from (Li et al., 2024).
Methods Training Paradigm Parameters Spearman $\uparrow$ Pearson $\uparrow$
Discriminative Model Baselines (Cross-Encoder Architecture)
[…]
Generative Large Language Model Baselines
[…]
Flan-UL2 Few-shot 20B 23.5 -
Tk-Instruct Few-shot 11B 17.8 -
GPT-3.5 Few-shot 175B 15.3 -
GPT-4 Few-shot - 43.6 -
Our Implementation on Qwen3-8B
Qwen3-8B Few-shot 8B 37.9 38.54
Qwen3-8B SFT 8B 40.42 40.83
PoLi-RL (Ours) RL 8B 48.18 48.27
3.2 Main Results
Table 1 summarizes the performance of our framework. PoLi-RL achieves Spearman and Pearson correlation coefficients of 48.18 and 48.27, respectively, establishing a new state-of-the-art (SOTA) for the cross-encoder architecture.
Comparing against prior work and different training paradigms reveals several key insights. First, PoLi-RL surpasses the previous cross-encoder SOTA, SEAVER (Li et al., 2024), by a significant margin of 4.35 points. Second, when compared to our own implementations on the same Qwen3-8B model, PoLi-RL achieves substantial improvements of 7.76 points over SFT and 10.28 points over few-shot prompting. The advantage over SFT is particularly noteworthy: while supervised fine-tuning relies on token-level cross-entropy loss that only implicitly encourages correct outputs, reinforcement learning enables explicit optimization through carefully designed reward signals that can directly guide the model’s reasoning process and better capture the nuanced requirements of conditional similarity judgment.
[…]
3.3 Ablation Studies
To understand the source of PoLi-RL’s performance gains and validate our design choices, we conduct a series of ablation studies.
Effectiveness of the Two-Stage Curriculum and Reward Components.
Table 2 dissects the effectiveness of our progressive training schedule. We first observe that a Naive RL approach (Row 2), which uses only a single, batch-wise listwise reward from scratch, yields negligible improvement over the few-shot baseline (Row 1), demonstrating the need for a more structured curriculum. Our PoLi-RL Stage I (Row 3) addresses this by building a robust foundation, substantially outperforming the few-shot baseline by 6.87 points. Within this stage, ablating the binary reward (Row 4) leads to a discernible dip in performance, validating its role in grounding the model in the task’s basic binary judgment.
Building upon this, the full PoLi-RL model (Row 5) further boosts performance by another 3.41 points (Row 5). Deconstructing the success of this final stage reveals that both ranking signals are vital: removing the listwise reward (Row 6) incurs the most significant penalty, while removing the pairwise reward (Row 7) also hinders performance. These findings confirm that both the two-stage curriculum and each of its constituent reward signals are essential for achieving optimal results, with the listwise signal being the most critical component in the final refinement stage.
Table 2: Ablation study on PoLi-RL’s two-stage training design and its reward components. The $\Delta$ column shows the absolute improvement in Spearman correlation over the indicated baseline.
Method Reward Component(s) Spearman $\uparrow$ Pearson $\uparrow$ $\Delta$ (Spearman)
(1) Few-shot Inference - 37.9 38.54 -
(2) Naive RL Listwise 38.19 38.34 +0.29 vs. (1)
(3) PoLi-RL (Stage I) Pointwise + Binary 44.77 44.45 +6.87 vs. (1)
(4)  - w/o Binary Pointwise 44.19 43.54 -0.58 vs. (3)
(5) PoLi-RL (Full) Pointwise + Pairwise + Listwise 48.18 48.27 +3.41 vs. (3)
(6)  - w/o Listwise Pointwise + Pairwise 46.71 46.37 -1.47 vs. (5)
(7)  - w/o Pairwise Pointwise + Listwise 47.6 47.59 -0.58 vs. (5)
Sensitivity to Reward Weights in PoLi-RL Stage II.
Table 3 details our analysis of the framework’s sensitivity to reward weights in Stage II. The results reveal a delicate balance, with peak performance achieved by moderately increasing the pairwise weight to 1.5. The model also shows considerable tolerance to pairwise signal, as halving its weight to 0.5 results in only a marginal performance drop to 47.77. This suggests that the stable, local ranking signal from the pairwise reward is instrumental in refining the model’s performance to its optimal state. In contrast, the framework is more sensitive to the pointwise and listwise weights, where deviations from their baseline of 1.0 tend to hinder performance. Crucially, the framework’s performance remains robust across all reward weight configurations. This demonstrates that our hybrid reward, anchored by the pointwise signal and refined by the PSRR mechanism, provides a consistently effective learning signal.
Table 3: Ablation study on the reward weights ($\mu_{1},\mu_{2},\mu_{3}$) in PoLi-RL’s Stage II.
Method $\mu_{1}$ (Pointwise) $\mu_{2}$ (Pairwise) $\mu_{3}$ (Listwise) Spearman $\uparrow$ Pearson $\uparrow$
PoLi-RL (Stage II) 1.0 1.0 1.0 47.83 47.83
1.0 1.5 1.0 48.18 48.27
1.5 1.0 1.0 47.3 47.23
1.0 1.0 1.5 47.46 47.48
1.0 0.5 1.0 47.77 47.31
0.5 1.0 1.0 47.36 47.18
1.0 1.0 0.5 47.39 47.27
Sensitivity to Parallel-Slice Size.
To determine the optimal configuration for our PSRR mechanism, we study the impact of the slice size $N$, with results presented in Table 4. The empirical results reveal a clear trend: performance peaks at an intermediate slice size of $N=24$ and degrades as the size deviates in either direction. This suggests that an optimal balance is required for the ranking signal. A slice that is too small may provide a less stable ranking signal, while one that is too large makes the ranking task overly complex for the model to learn effectively. This finding validates the design principle behind our PSRR mechanism: a carefully-sized, localized ranking signal is more effective than a purely global or an overly-restricted one.
Table 4: Analysis on the impact of the parallel slice size (NN) in Stage II. NN represents the number of samples used for listwise ranking reward computation in each slice.
Method NN (Slice Size) Spearman $\uparrow$ Pearson $\uparrow$
PoLi-RL(Stage II) 16 47.16 46.96
24 48.18 48.27
32 47.44 47.19
40 47.18 47.32
48 46.78 46.84
4 Analysis
In this section, we provide a quantitative and qualitative analysis of PoLi-RL’s performance, focusing on its error distribution and nuanced reasoning capabilities.
4.1 Analysis of Prediction Error Distribution
Figure 3 visualizes the distribution of absolute prediction errors on the C-STS validation set. The plot reveals several key insights into the models’ behaviors. While an error of 1 is the most frequent outcome for all models, likely reflecting the inherent nuances of the C-STS scale, a clear progression of improvement is evident. Compared to the raw and SFT model, PoLi-RL demonstrates a more favorable error distribution. Firstly, it achieves the highest density of perfect predictions (error=0). More importantly, the density of the PoLi-RL curve in the high-error regions is consistently the lowest. This shows that our method significantly reduces the frequency of large, unreliable errors, yielding a more stable and reliable model.
Figure 3: Distribution of absolute prediction errors for the raw model, SFT model, and PoLi-RL. The $x$-axis represents the absolute error ($\left|\text{Predicted}-\text{True}\right|$), and the $y$-axis is the probability density.
4.2 Qualitative Analysis: A Case Study on Nuanced Reasoning
To qualitatively analyze our framework’s nuanced reasoning ability, we present a case study in Figure 4 on the challenging condition “The person’s connection with the ground”. The baseline models show clear deficiencies: the Few-shot model exhibits a brittle, literal interpretation, fixating on the preposition ‘on’ while ignoring the real-world context, leading to an incorrect prediction. The SFT Baseline, while achieving a closer score, shows uncertainty and fails to make the necessary fine-grained distinction. In contrast, PoLi-RL applies world knowledge to infer that both “jumping” and “doing a flip” result in the same “not connected” state, arriving at the accurate score of 4 by identifying the differing actions leading to the state as unimportant details. This case study demonstrates that PoLi-RL cultivates the contextual inference ability required for complex conditional judgments.
Case Study: Model Comparison on a Reasoning Sample Sentence 1: A person riding a snowboard jumps high over the snowy hill. Sentence 2: A man in a blue snowsuit does a flip on a huge snowy mountain. Condition: The person’s connection with the ground. Gold Label: 4 (Mostly equivalent) Few-shot Inference Prediction: no(1) Reasoning: In Sentence 1… , the key action is jumping, which implies the person is airborne and not in contact with the ground. In Sentence 2… , the preposition ‘on’ suggest that he’s on the mountain’s surface. So, he’s connected to the ground. Therefore, the two sentences have different connections with the ground. Hence, the answer is ‘no(1)’. SFT Baseline Prediction: yes(3) PoLi-RL Prediction: yes(4) Reasoning: Sentence 1 describes… . So he’s airborne during the jump, meaning he’s not connected to the ground. Sentence 2 talks about … I infer that this action, similar to a jump or a trick in a snow-sports context, also means he is in the air. Therefore, they are similar because both have the person not connected to the ground. But the final score should be 4 because the action (jumping/flipping) leading to this state represent minor, differing details.
Figure 4: A detailed case study comparing the reasoning processes and final outputs of the Few-shot, SFT baselines, and PoLi-RL on a C-STS sample that requires nuanced reasoning.
5 Conclusion
We introduce PoLi-RL, a two-stage reinforcement learning framework that resolves the complex credit assignment problem associated with listwise objectives in C-STS through a progressive curriculum and an innovative Parallel-Slice Ranking Reward (PSRR) mechanism. Our method establishes a new state-of-the-art for the cross-encoder architecture, significantly outperforming SFT baselines and even large models like GPT-4. As the first successful application of reinforcement learning to this task, our study validates a powerful paradigm for aligning LLMs with complex ranking objectives, paving the way for future applications in retrieval and other ranking-based tasks.
6 Contact Information
Emails:    Zixin Song: songzx24@mails.tsinghua.edu.cn    Bowen Zhang: zbw23@mails.tsinghua.edu.cn
Qian-Wen Zhang: cowenzhang@tencent.com
Appendix A Appendix
A.1 Related Work
Conditional Semantic Textual Similarity
C-STS is a recent advancement over traditional STS that introduces a natural language condition to disambiguate the measurement of similarity between two texts. Research in this area has primarily established three mainstream paradigms: Bi-encoder, Tri-encoder, and Cross-encoder. Given two texts and a condition, the bi-encoder architecture typically uses a Siamese network to encode two text-condition pairs, while the tri-encoder architecture encodes each text and the condition separately before an aggregation step. A prevalent optimization strategy for these paradigms is contrastive learning. For instance, [4] propose a conditional contrastive learning framework that pulls representations of the same text pair under a high-similarity condition closer, while pushing them apart under a low-similarity one. Extending this, Hyper-CL [20] utilizes a tri-encoder setup where a hypernetwork generates condition-specific projectors to dynamically adapt sentence representations within a contrastive objective. More recently, CSR introduced a parameter-free router for the tri-encoder, using the condition to re-weight sentence tokens to amplify relevant information without increasing model size.
In contrast, the cross-encoder architecture processes the concatenated texts and condition as a single input, enabling deep, token-level interaction. However, this theoretical advantage did not consistently translate to superior performance in earlier discriminative models. The state-of-the-art method in this setting, SEAVER, addressed this discrepancy by identifying that such models can be distracted by condition-irrelevant tokens. To resolve this, SEAVER [6] introduces an attention reallocation mechanism that optimizes the model by re-weighting internal attention scores during fine-tuning, forcing a focus on the most salient information.
The advent of LLMs has introduced new approaches, mainly few-shot inference and using LLMs as feature extractors [7] [21]. However, all these prior works, regardless of architecture, are confined to supervised paradigms like SFT or contrastive learning. Our work is the first to apply reinforcement learning to this task.
Reinforcement Learning for Large Language Models
Reinforcement Learning (RL) is pivotal for aligning and enhancing Large Language Models (LLMs), with algorithms like Proximal Policy Optimization (PPO) [22] widely used to optimize non-differentiable objectives in tasks such as reasoning and code generation. To address PPO’s limitations, such as high variance in long-sequence tasks, advanced variants have emerged.
Group Relative Policy Optimization (GRPO) [8] addresses these limitations by introducing a group-relative advantage estimation, which eliminates the need for a separate value function through Z-score normalization of rewards within sample groups. This approach simplifies training and enhances sample efficiency, as demonstrated in models like DeepSeek-R1. Building on this, Decoupled Clip and Dynamic Sampling Policy Optimization (DAPO) [13] provides an open-source, scalable RL system tailored for LLMs. DAPO incorporates key improvements, including dynamic sampling to adaptively adjust the number of generated completions based on reward variance, making it particularly effective for long-horizon reasoning tasks.
Leveraging DAPO’s powerful optimization engine, our work, PoLi-RL, marks a significant departure by being the first to introduce a reinforcement learning framework to the C-STS task, thereby establishing a new optimization paradigm.
A.2 Prompt Template for PoLi-RL
Below is the detailed few-shot prompt used for both the few-shot inference baseline and the training, evaluation process of PoLi-RL.
Prompt for C-STS Task Judge the semantic similarity between Sentence 1 and Sentence 2 based completely on the given Condition. The final output must be exactly in this format: the similarity judgment (‘yes’ or ‘no’) followed by the score in parentheses, wrapped in <answer></answer> tags. Examples: <answer>yes(4)</answer>, <answer>no(1)</answer>. Include no other text, tags, or explanations. To arrive at this output, follow these two steps: Step 1: Binary Judgment. Determine if the sentences are ‘similar’ (‘yes’) or ‘not similar’ (‘no’). ‘similar’: The sentences are roughly, mostly, or completely equivalent under the condition. ‘not similar’: The sentences are dissimilar under the condition. Step 2: Fine-grained Score. Assign an integer score based on Step 1: For a ‘yes’ judgment: * 5: The two sentences are completely equivalent as they mean the same thing with respect to the condition. * 4: The two sentences are mostly equivalent, but some unimportant details differ with respect to the condition. * 3: The two sentences are roughly equivalent, but some important information differs or is missing with respect to the condition. For a ‘no’ judgment: * 2: The two sentences are dissimilar, but are on a similar topic with respect to the condition or shares a close semantic relationship. This applies when items are clearly different, but not direct opposites. * 1: The two sentences are dissimilar with respect to the condition, representing a direct opposition or a clear, unrelated difference. (e.g., ‘man’ vs. ‘woman’). Here are some examples: Example 1:
<Sentence1>: A girl is cooking in a kitchen and a man is standing next to her.
<Sentence2>: A man sitting with a pizza in his hand in front of pizza on the table.
<Condition>: The number of people.
<answer>no(1)</answer>
Explanation: The first sentence mentions two people, while the second sentence mentions only one person.
Example 2:
<Sentence1>: A wood table sitting by a wood framed bed with a lamp on it.
<Sentence2>: A microwave, refrigerator, television, and wooden drawers sit in the corner of a bedroom.
<Condition>: The room type.
<answer>yes(5)</answer>
Explanation: We can infer from the two sentences that the room type are both bedroom.
Example 3:
<Sentence1>: A small crowd gathered around the injured person.
<Sentence2>: A crowd jumps up and down to the tunes played by an artist.
<Condition>: The number of people
<answer>yes(3)</answer>
Explanation: While both sentences mention crowds, it is important and unclear how many people there are.
Now, apply these steps to the following sentences: <Sentence1>: {sentence1}
<Sentence2>: {sentence2}
<Condition>: {condition}